# Exact pp-wave solutions in shift-symmetric higher-order scalar-tensor theories

### 19 Mar 2026 v1

@@ Abstract
We investigate exact plane-fronted gravitational wave (pp-wave) solutions within the framework of shift-symmetric quadratic-order higher-order scalar–tensor (HOST) theories. These solutions represent fully nonlinear radiative spacetimes that extend beyond the linearized approximation. We demonstrate that under the algebraic conditions on the coupling functions, the gravitational field equations reduce to a two-dimensional Laplace equation for the wave profile, recovering the structural form of vacuum general relativity (GR). By adopting a scalar field ansatz that depends linearly on transverse coordinates and arbitrarily on the retarded null coordinate, we maintain a constant kinetic term of the scalar field. This configuration allows for a stealth pp-wave solution, where a nontrivial scalar field profile coexists with the gravitational wave without backreacting on the spacetime geometry. We further show that these stealth configurations are fully compatible with the degeneracy conditions of Class-Ia DHOST theories and satisfy current observational constraints. Finally, we examine the behavior of these solutions under disformal transformations, revealing that while the Brinkmann form is preserved, the stealth property is generically lost due to the mixing of scalar and tensor degrees of freedom. These results establish the robustness of pp-wave solutions in viable DHOST frameworks and highlight their utility for probing nonlinear effects in modified gravity.
@@ I Introduction
From the theoretical viewpoint, scalar–tensor extensions of GR have received considerable attention as candidate theories of dark energy and modified gravity [7, 8, 9, 10, 11, 12]. In particular, degenerate higher-order scalar–tensor (DHOST) theories provide the most general class of scalar–tensor theories with higher-derivative interactions that remain free of Ostrogradsky instabilities [13, 14, 15]. These theories cover a wide variety of models, including Horndeski and beyond-Horndeski theories [8, 9, 10, 11], and are characterized by a set of degeneracy conditions ensuring the absence of extra ghostly degrees of freedom. Observational constraints from the coincidence of the speed of gravitational waves and light, inferred from the binary neutron star merger GW170817 [16, 17, 18, 19], further restrict viable subclasses of DHOST theories [20, 21, 22, 23, 24]. Furthermore, no perturbative or nonperturbative decay of gravitational waves into dark energy imposes further constraints on the theories [25, 26].
@@
We would like to ask whether the exact pp-wave solutions of GR continue to exist within the framework of DHOST theories. In this work, we focus on quadratic-order shift-symmetric higher-order scalar–tensor (HOST) theories without imposing the degeneracy conditions. Such theories are constructed from quadratic combinations of the second-order derivatives of the scalar field, and include in addition generalized kinetic term, generalized cubic Galileon term, and the nonminimal coupling to the Ricci tensor. We first examine whether the pp-wave solutions remain valid solutions within quadratic-order shift-symmetric HOST theories. We then investigate their compatibility with the degeneracy conditions that characterize Class Ia quadratic-order DHOST theories, as well as with additional constraints arising from observational viability, such as the consistency with gravitational-wave propagation.
@@
Starting from a generalized pp-wave ansatz for the metric and a scalar field profile depending on the null and transverse coordinates, we derive the corresponding Euler–Lagrange equations and then investigate whether these Euler-Lagrange equations are satisfied for the pp-wave metric and the form of the scalar field compatible with the shift symmetry. If such solutions exist, they could be useful for investigating the propagation of gravitational and scalar waves in modified gravity, and clarify the interplay between nonlinear gravitational radiation and scalar degrees of freedom. The possible existence of stealth scalar configurations, which are nontrivial scalar field profiles that do not backreact on the background geometry and are originally investigated for static or stationary black hole solutions [27, 28, 29, 30, 31, 32, 33, 34], offers an intriguing mechanism by which scalar degrees of freedom may coexist with gravitational waves without altering their spacetime structure.
@@
Previous studies of gravitational waves in modified gravity theories have mostly focused on the linearized regime, considering small perturbations around Minkowski or cosmological backgrounds [35]. By contrast, the fate of exact nonlinear radiative solution such as pp-wave configurations have not been unexplored much in higher-order scalar–tensor (HOST) theories, unless one performs a fully nonlinear analysis. Understanding these solutions is crucial for assessing the robustness of gravitational-wave propagation and their potential interactions with additional scalar degrees of freedom inherent in modified gravity frameworks.
@@
Exact solutions provide genuinely nonperturbative information about the dynamics of the theory that cannot be captured by linearized analyses, revealing effects that may only appear at strong-field or highly nonlinear regimes. Such solutions serve as theoretical laboratories for systematically testing fundamental aspects of the propagation of scalar and tensor degrees of freedom, their mutual interactions, and the possible existence of stealth configurations where scalar fields coexist with gravitational waves without backreacting on the spacetime geometry. They help to clarify the consistency of the theory with the fundamental principle that gravitational radiation should propagate without pathological behavior even in the high-frequency limit or in the presence of strong-field effects.
@@
We will find that under only a few mild conditions on the coupling functions the resulting Euler–Lagrange equations reduce to a two-dimensional Laplace equation for the wave profile $H(u,x_{1},x_{2})$, precisely matching the condition required for the existence of exact pp-wave solutions in vacuum GR. The scalar field is then chosen to have a linear dependence on the transverse coordinates, supplemented by an arbitrary function of the retarded time $u$, which leads to a constant kinetic term $X=X_{0}$, where $X$ is the kinetic term of the scalar field (see II), and is fully compatible with the shift symmetry of the theory. Imposing consistency with the Euler–Lagrange equations constrains only simple algebraic conditions on the potential function $F_{0}(X)$ evaluated at $X=X_{0}$. When these conditions are satisfied, the scalar field propagates alongside the gravitational wave without affecting the spacetime geometry—a configuration we interpret as a stealth pp-wave solution. We further demonstrate that such pp-wave solutions of GR persist in the viable subclasses of quadratic-order shift-symmetric DHOST theories, accompanied by a scalar field that does not backreact on the metric. This result demonstrates the robustness of pp-waves as exact solutions beyond GR, reinforces their utility as theoretical probes for exploring modified gravity, and opens a pathway for future studies of nonlinear gravitational radiation in higher-order scalar–tensor frameworks, including potential applications to observational signatures and fundamental consistency tests.
@@ II The model
We consider the quadratic-order Higher-Order Scalar-Tensor (HOST) theories introduced in [13], whose Lagrangian contains quadratic contractions of the second derivatives of the scalar field. The action of the HOST theories is given by
@@
where $\phi_{\mu\nu}:=\nabla_{\nu}\nabla_{\mu}\phi$ denotes the second covariant derivative of $\phi$, and $\Box\phi:=g^{\mu\nu}\phi_{\mu\nu}$ is the d’Alembertian. The functions $F_{i}(\phi,X)$ ($i=0,1,2$) and $A_{I}(\phi,X)$ ($I=1,\ldots,5$) are arbitrary scalar functions of their arguments, thereby parameterizing a very general class of higher-order scalar-tensor interactions.
The quadratic-order Degenerate Higher-Order Scalar-Tensor (DHOST) theories arise as special subclasses of Eq. (1), identified by degeneracy conditions that eliminate unwanted Ostrogradsky instabilities. Within this large family, a particularly relevant class is provided by the Class-Ia DHOST theories, which ensure that the propagation speed of tensor modes ($c_{t}$) matches the speed of light ($c$). Such theories are motivated by the stringent observational bounds on the speed of gravitational waves from the binary neutron star merger GW170817 [23]. The degeneracy conditions specifying this subclass read
@@
where $A_{3}$ remains a free function of $(\phi,X)$. An additional physical requirement further constrains this subclass. In particular, it has been argued that in the presence of a non-vanishing $A_{3}$, gravitational waves would generically suffer a perturbative or even non-perturbative decay into fluctuations of the scalar sector, interpreted as dark energy [25, 26]. To forbid such an instability, one must impose
@@
Finally, we impose a global shift symmetry $\phi\to\phi+c$, with $c$ constant. Under this assumption, all functions in the action (1) depend solely on the kinetic term $X$, thereby eliminating explicit $\phi$-dependence. The shift symmetry not only reduces the parameter space of admissible theories but also plays an important role in protecting the scalar sector from large quantum corrections. Consequently, the setup described above represents one of the most theoretically consistent and observationally viable corners of the quadratic HOST/DHOST landscape.
III Plane-fronted gravitational wave solutions
@@
where $u$ is the retarded null coordinate, representing the phase of the wave, and $v$ is the affine parameter along the null geodesics associated with the covariantly constant null vector field $\ell^{\mu}=\left(\partial_{v}\right)^{\mu}$. It should be noted that $v$ is not a null coordinate, since the hypersurfaces $v=\text{const}$ are not null. $(x_{1},x_{2})$ are Cartesian coordinates on the two-dimensional transverse space orthogonal to the propagation direction. $H(u,x_{1},x_{2})$ is an arbitrary function specifying the wave profile.
and $(x_{1},x_{2})$ are cartesian coordinates on the two-dimensional transverse space orthogonal to the propagation direction. $H(u,x_{1},x_{2})$ is an arbitrary function specifying the wave profile. The wave propagates in the direction of increasing $u$, with each surface $u=\mathrm{const}$ representing a wavefront of constant phase. These are null hypersurfaces that carry the entire curvature content of the spacetime. The pp-wave spacetimes are members of the Kundt class of metrics and are characterized by the existence of a covariantly constant null vector field $\ell^{\mu}=(\partial_{v})^{\mu}$, satisfying
@@
Any $H(u,x_{1},x_{2})$ satisfying Eq. (9) yields an exact Ricci-flat pp-wave spacetime. We note that Eq. (9) is linear in $x_{1}$ and $x_{2}$, even though the underlying Einstein equations in GR are fully nonlinear in the metric components. This means that the transverse spatial profile of the wave can be obtained by superposition of harmonic functions, but the dependence on $u$ can be completely arbitrary, allowing for highly nontrivial temporal behavior. A particularly simple and physically relevant subclass of pp-wave metrics is obtained by choosing $H$ to be quadratic in the transverse coordinates:
@@
This clearly shows that $A(u)$ and $B(u)$ govern the tidal stretching and squeezing experienced by particles in the transverse plane, providing a direct operational meaning to the gravitational wave polarizations in the exact theory.
Conserved quantities in radiative spacetimes are naturally defined at future null infinity within the Bondi–Sachs framework, where the Bondi mass decreases according to the standard mass-loss formula involving the news tensor $N_{AB}$. Pure pp-wave geometries, however, are not asymptotically flat in the spherical sense and instead admit a planar-type null infinity generated by the null Killing vector $\ell^{\mu}=\partial_{v}$. In this setting the Bondi-type mass vanishes identically, while the analogue of the news tensor in the transverse plane is nonzero whenever $\partial_{u}H\neq 0$, reflecting the radiative nature of the wave. Although the mass is zero, conserved momentum and angular-momentum fluxes arise from Killing symmetries of the transverse plane, and in vacuum pp-waves these quantities are purely geometrical.
@@
In order to investigate whether the vacuum pp-wave solutions of GR also persist as exact solutions within the framework of HOST theories, we begin with the most general metric ansatz compatible with the existence of a null coordinate and a two-dimensional Euclidean transverse space. From a purely geometric perspective, pp-waves are characterized by the existence of a covariantly constant null vector field $\ell$ satisfying Eq. (7), i.e., $\nabla_{\nu}\ell_{\mu}=0$. This condition indicates, via the Frobenius theorem, that $\ell_{\mu}$ is hypersurface orthogonal ($\ell_{[\mu}\nabla_{\nu}\ell_{\rho]}=0$), which allows us to label the corresponding null hypersurfaces with a null coordinate $u$. Furthermore, because $\ell$ is geodesic by definition, we can choose the affine parameter $v$ along the null rays as a second coordinate such that $\ell^{\mu}=(\partial_{v})^{\mu}$. Combined with the two transverse coordinates $(x_{1},x_{2})$ of a Euclidean 2-space, this geometric structure strongly restricts the form of the metric and its coordinate dependence from the outset. Importantly, we do not assume that a pp-wave solves the HOST field equations; the restrictions imposed below follow solely from coordinate freedom and symmetry considerations.
@@
See Appendix A for details of the gauge fixing. Since these simplifications rely solely on coordinate freedom, and not on the field equations, the metric can always be written, without loss of generality, as
@@
In the following, we investigate whether this specific geometric class admits solutions within the framework of HOST theories. After deriving the Euler–Lagrange equation for the scalar field, we adopt the ansatz
@@
The remaining Euler–Lagrange equations reduce to the algebraic compatibility conditions
@@
The resulting solution represents a scalar field that propagates together with a gravitational pp-wave while remaining completely hidden from the gravitational dynamics. The term “stealth” refers to the remarkable situation in which the scalar field is fully nontrivial, with both a wave component depending on the null coordinate $u$ and constant gradients in the transverse directions, yet its energy–momentum tensor vanishes once the field equations are imposed. This vanishing is not a consequence of the scalar field being trivial. It arises from a delicate cancellation among the various contributions that appear in HOST theories, including terms involving first and second derivatives of the scalar field. The cancellation ensures that the scalar field does not act as a source for the geometry even though it carries nonzero gradients and possesses genuine dynamical content.
Because the stress tensor disappears on-shell, the geometry is identical to that of a vacuum pp-wave in general relativity. The curvature is determined entirely by the harmonic function $H(u,x_{1},x_{2})$, and the characteristic structure of pp-waves is preserved. The null vector $\partial_{v}$ remains covariantly constant, the Weyl tensor keeps its type N form, and the tidal forces experienced by test particles are exactly those of a pure gravitational wave. No observable quantity in the gravitational sector reveals the presence of the scalar field. From the viewpoint of GR alone, the spacetime would be interpreted as a standard vacuum pp-wave without any additional degrees of freedom.

### 31 Mar 2026 v2

@@ Abstract
We investigate exact plane-fronted gravitational-wave (pp-wave) solutions within the framework of shift-symmetric quadratic-order higher-order scalar-tensor (HOST) theories. These solutions represent fully nonlinear radiative spacetimes that extend beyond the linearized approximation. We demonstrate that under the algebraic conditions on the coupling functions, the gravitational field equations reduce to a two-dimensional Laplace equation for the wave profile, recovering the structural form of vacuum general relativity. By adopting a scalar field ansatz that depends linearly on transverse coordinates and arbitrarily on the retarded null coordinate, we maintain a constant kinetic term of the scalar field. This configuration allows for a “stealth pp-wave” solution, where a nontrivial scalar field profile coexists with the gravitational wave without backreacting on the spacetime geometry. We further show that these stealth configurations are fully compatible with the degeneracy conditions of class-Ia DHOST theories and satisfy current observational constraints. Finally, we examine the behavior of these solutions under disformal transformations, revealing that while the Brinkmann form is preserved, the stealth property is generically lost due to the mixing of scalar and tensor degrees of freedom. These results establish the robustness of pp-wave solutions in viable DHOST frameworks and highlight their utility for probing nonlinear effects in modified gravity.
@@ I Introduction
From the theoretical viewpoint, scalar-tensor extensions of GR have received considerable attention as candidate theories of dark energy and modified gravity [7, 8, 9, 10, 11, 12]. In particular, degenerate higher-order scalar–tensor (DHOST) theories provide the most general class of scalar-tensor theories with higher-derivative interactions that remain free of Ostrogradsky instabilities [13, 14, 15]. These theories cover a wide variety of models, including Horndeski and beyond-Horndeski theories [8, 9, 10, 11], and are characterized by a set of degeneracy conditions ensuring the absence of extra ghostly degrees of freedom. Observational constraints from the coincidence of the speed of gravitational waves and light, inferred from the binary neutron star merger GW170817 [16, 17, 18, 19], further restrict viable subclasses of DHOST theories [20, 21, 22, 23, 24]. Furthermore, no perturbative or nonperturbative decay of gravitational waves into dark energy imposes further constraints on the theories [25, 26].
@@
We would like to ask whether the exact pp-wave solutions of GR continue to exist within the framework of DHOST theories. In this work, we focus on quadratic-order shift-symmetric higher-order scalar-tensor (HOST) theories without imposing the degeneracy conditions. Such theories are constructed from quadratic combinations of the second-order derivatives of the scalar field, and include, in addition, a generalized kinetic term, a generalized cubic Galileon term, and the nonminimal coupling to the Ricci tensor. We first examine whether the pp-wave solutions remain valid solutions within quadratic-order shift-symmetric HOST theories. We then investigate their compatibility with the degeneracy conditions that characterize class Ia quadratic-order DHOST theories, as well as with additional constraints arising from observational viability, such as the consistency with gravitational-wave propagation.
@@
Starting from a generalized pp-wave ansatz for the metric and a scalar field profile depending on the null and transverse coordinates, we derive the corresponding Euler-Lagrange equations and then investigate whether these Euler-Lagrange equations are satisfied for the pp-wave metric and the form of the scalar field compatible with the shift symmetry. If such solutions exist, they could be useful for investigating the propagation of gravitational and scalar waves in modified gravity, and clarify the interplay between nonlinear gravitational radiation and scalar degrees of freedom. The possible existence of stealth scalar configurations, which are nontrivial scalar field profiles that do not backreact on the background geometry and are originally investigated for static or stationary black hole solutions [27, 28, 29, 30, 31, 32, 33, 34], offers an intriguing mechanism by which scalar degrees of freedom may coexist with gravitational waves without altering their spacetime structure.
@@
Previous studies of gravitational waves in modified gravity theories have mostly focused on the linearized regime, considering small perturbations around Minkowski or cosmological backgrounds [35]. By contrast, the fate of exact nonlinear radiative solution, such as pp-wave configurations, have not been unexplored much in HOST theories, unless one performs a fully nonlinear analysis. Understanding these solutions is crucial for assessing the robustness of gravitational-wave propagation and their potential interactions with additional scalar degrees of freedom inherent in modified gravity frameworks.
@@
Exact solutions provide genuinely nonperturbative information about the dynamics of the theory that cannot be captured by linearized analyses, revealing effects that may only appear at strong-field or highly nonlinear regimes [36, 37, 38]. Such solutions serve as theoretical laboratories for systematically testing fundamental aspects of the propagation of scalar and tensor degrees of freedom, their mutual interactions, and the possible existence of stealth configurations where scalar fields coexist with gravitational waves without backreacting on the spacetime geometry. They help to clarify the consistency of the theory with the fundamental principle that gravitational radiation should propagate without pathological behavior even in the high-frequency limit or in the presence of strong-field effects.
@@
We will find that under only a few mild conditions on the coupling functions the resulting Euler-Lagrange equations reduce to a two-dimensional Laplace equation for the wave profile $H(u,x_{1},x_{2})$, precisely matching the condition required for the existence of exact pp-wave solutions in vacuum GR. The scalar field is then chosen to have a linear dependence on the transverse coordinates, supplemented by an arbitrary function of the retarded time $u$, which leads to a constant kinetic term $X=X_{0}$, where $X$ is the kinetic term of the scalar field (see Sec. II), and is fully compatible with the shift symmetry of the theory. Imposing consistency with the Euler-Lagrange equations constrains only simple algebraic conditions on the potential function $F_{0}(X)$ evaluated at $X=X_{0}$. When these conditions are satisfied, the scalar field propagates alongside the gravitational wave without affecting the spacetime geometry—a configuration we interpret as a stealth pp-wave solution. We further demonstrate that such pp-wave solutions of GR persist in the viable subclasses of quadratic-order shift-symmetric DHOST theories, accompanied by a scalar field that does not backreact on the metric. This result demonstrates the robustness of pp-waves as exact solutions beyond GR, reinforces their utility as theoretical probes for exploring modified gravity, and opens a pathway for future studies of nonlinear gravitational radiation in higher-order scalar-tensor frameworks, including potential applications to observational signatures and fundamental consistency tests.
@@ II The model
We consider the quadratic-order HOST theories introduced in [13], whose Lagrangian contains quadratic contractions of the second derivatives of the scalar field. The action of the HOST theories is given by
@@
where $\phi_{\mu\nu}:=\nabla_{\nu}\nabla_{\mu}\phi$ denotes the second covariant derivative of $\phi$, and $\Box\phi:=g^{\mu\nu}\phi_{\mu\nu}$ is the d’Alembertian. The functions $F_{i}(\phi,X)$ ($i=0,1,2$) and $A_{I}(\phi,X)$ ($I=1,\ldots,5$) are arbitrary scalar functions of their arguments, thereby parametrizing a very general class of higher-order scalar-tensor interactions.
The quadratic-order DHOST theories arise as special subclasses of Eq. (1), identified by degeneracy conditions that eliminate unwanted Ostrogradsky instabilities. Within this large family, a particularly relevant class is provided by the class-Ia DHOST theories, which ensure that the propagation speed of tensor modes ($c_{t}$) matches the speed of light ($c$). Such theories are motivated by the stringent observational bounds on the speed of gravitational waves from the binary neutron star merger GW170817 [23]. The degeneracy conditions specifying this subclass read
@@
where $A_{3}$ remains a free function of $(\phi,X)$. An additional physical requirement further constrains this subclass. In particular, it has been argued that, in the presence of a nonvanishing $A_{3}$, gravitational waves would generically suffer a perturbative or even nonperturbative decay into fluctuations of the scalar sector, interpreted as dark energy [25, 26]. To forbid such an instability, one must impose
@@
Finally, we impose a global shift symmetry $\phi\to\phi+c$, with $c$ constant. Under this assumption, all functions in the action (1) depend solely on the kinetic term $X$, thereby eliminating explicit $\phi$ dependence. The shift symmetry not only reduces the parameter space of admissible theories but also plays an important role in protecting the scalar sector from large quantum corrections. Consequently, the setup described above represents one of the most theoretically consistent and observationally viable corners of the quadratic HOST/DHOST landscape.
III Plane-fronted gravitational-wave solutions
@@
where $u$ is the retarded null coordinate, representing the phase of the wave, and $v$ is the affine parameter along the null geodesics associated with the covariantly constant null vector field $\ell^{\mu}=\left(\partial_{v}\right)^{\mu}$. It should be noted that $v$ is not a null coordinate, since the hypersurfaces $v=\text{const}$ are not null. $(x_{1},x_{2})$ are Cartesian coordinates on the two-dimensional transverse space orthogonal to the propagation direction. $H(u,x_{1},x_{2})$ is an arbitrary function specifying the wave profile. The wave propagates in the direction of increasing $u$, with each surface $u=\mathrm{const}$ representing a wavefront of constant phase. These are null hypersurfaces that carry the entire curvature content of the spacetime. The pp-wave spacetimes are members of the Kundt class of metrics and are characterized by the existence of a covariantly constant null vector field $\ell^{\mu}=(\partial_{v})^{\mu}$, satisfying
@@
Any $H(u,x_{1},x_{2})$ satisfying Eq. (9) yields an exact Ricci-flat pp-wave spacetime. We note that Eq. (9) is linear in $x_{1}$ and $x_{2}$, even though the underlying Einstein equations in GR are fully nonlinear in the metric components. This means that the transverse spatial profile of the wave can be obtained by superposition of harmonic functions, but the dependence on $u$ can be completely arbitrary, allowing for highly nontrivial temporal behavior. A particularly simple and physically relevant subclass of pp-wave metrics is obtained by choosing $H$ to be quadratic in the transverse coordinates,
@@
This clearly shows that $A(u)$ and $B(u)$ govern the tidal stretching and squeezing experienced by particles in the transverse plane, providing a direct operational meaning to the gravitational-wave polarizations in the exact theory.
Conserved quantities in radiative spacetimes are naturally defined at future null infinity within the Bondi-Sachs framework, where the Bondi mass decreases according to the standard mass-loss formula involving the news tensor $N_{AB}$. Pure pp-wave geometries, however, are not asymptotically flat in the spherical sense and instead admit a planar-type null infinity generated by the null Killing vector $\ell^{\mu}=\partial_{v}$. In this setting the Bondi-type mass vanishes identically, while the analog of the news tensor in the transverse plane is nonzero whenever $\partial_{u}H\neq 0$, reflecting the radiative nature of the wave. Although the mass is zero, conserved momentum and angular-momentum fluxes arise from Killing symmetries of the transverse plane, and in vacuum pp-waves these quantities are purely geometrical.
@@
In order to investigate whether the vacuum pp-wave solutions of GR also persist as exact solutions within the framework of HOST theories, we begin with the most general metric ansatz compatible with the existence of a null coordinate and a two-dimensional Euclidean transverse space. From a purely geometric perspective, pp-waves are characterized by the existence of a covariantly constant null vector field $\ell$ satisfying Eq. (7), i.e., $\nabla_{\nu}\ell_{\mu}=0$. This condition indicates, via the Frobenius theorem, that $\ell_{\mu}$ is hypersurface orthogonal ($\ell_{[\mu}\nabla_{\nu}\ell_{\rho]}=0$), which allows us to label the corresponding null hypersurfaces with a null coordinate $u$. Furthermore, because $\ell$ is geodesic by definition, we can choose the affine parameter $v$ along the null rays as a second coordinate such that $\ell^{\mu}=(\partial_{v})^{\mu}$. Combined with the two transverse coordinates $(x_{1},x_{2})$ of a Euclidean two-space, this geometric structure strongly restricts the form of the metric and its coordinate dependence from the outset. Importantly, we do not assume that a pp-wave solves the HOST field equations; the restrictions imposed below follow solely from coordinate freedom and symmetry considerations.
@@
See Appendix A for details of the gauge fixing. Since these simplifications rely solely on coordinate freedom and not on the field equations, the metric can always be written, without loss of generality, as
@@
In the following, we investigate whether this specific geometric class admits solutions within the framework of HOST theories. After deriving the Euler-Lagrange equation for the scalar field, we adopt the ansatz
@@
The remaining Euler-Lagrange equations reduce to the algebraic compatibility conditions
@@
The resulting solution represents a scalar field that propagates together with a gravitational pp-wave while remaining completely hidden from the gravitational dynamics. The term “stealth” refers to the remarkable situation in which the scalar field is fully nontrivial, with both a wave component depending on the null coordinate $u$ and constant gradients in the transverse directions, yet its energy-momentum tensor vanishes once the field equations are imposed. This vanishing is not a consequence of the scalar field being trivial. It arises from a delicate cancellation among the various contributions that appear in HOST theories, including terms involving first and second derivatives of the scalar field. The cancellation ensures that the scalar field does not act as a source for the geometry even though it carries nonzero gradients and possesses genuine dynamical content.
Because the stress tensor disappears on shell, the geometry is identical to that of a vacuum pp-wave in general relativity. The curvature is determined entirely by the harmonic function $H(u,x_{1},x_{2})$, and the characteristic structure of pp-waves is preserved. The null vector $\partial_{v}$ remains covariantly constant, the Weyl tensor keeps its type N form, and the tidal forces experienced by test particles are exactly those of a pure gravitational wave. No observable quantity in the gravitational sector reveals the presence of the scalar field. From the viewpoint of GR alone, the spacetime would be interpreted as a standard vacuum pp-wave without any additional degrees of freedom.